OpenStudy (anonymous):

If f(x) = 5x + 4 and g(x) = 6x – 3, what is f(g(x))?

OpenStudy (anonymous):

F(x) = 5x+4 G(x) = 6x-3 F(G(x))= F(6x-3)= 5(6x-3) +4 Which is equal to, 30x -15 +4 Or, 30x - 11 is the answer.

OpenStudy (nenadmatematika):

correct :D

OpenStudy (anonymous):

thank you:)

OpenStudy (nenadmatematika):

you're welcome yasmeena :D

OpenStudy (anonymous):

^^ lol only family calls me that?

OpenStudy (anonymous):

Haha no problem, btw Yasmeen try to learn the formulae its much easier xD

OpenStudy (nenadmatematika):

how do you want me to call you? :D

OpenStudy (anonymous):

lol nothing that was just funny

OpenStudy (nenadmatematika):

:D

OpenStudy (anonymous):

you dont mind me askin where u from?

OpenStudy (nenadmatematika):

from Belgrade, Serbia

OpenStudy (anonymous):

so your ethnicity is?

OpenStudy (nenadmatematika):

I'm Serbian